OpenStudy (anonymous):

what is the area of the circle whose equation is (x+1)^2+(y+1)^2=1?

OpenStudy (anonymous):

round your answer to the nearest hundredth.

OpenStudy (goformit100):

@satellite73

OpenStudy (anonymous):

radius of \((x-h)^2+(y-k)^2=r^2\) is \(r\) and the area of a circle with radius \(r\) is \(\pi r^2\)

OpenStudy (anonymous):

in your case \(r^2=1\) and so \(r=1\)

OpenStudy (anonymous):

thank you ! @satellite73 @goformit100

OpenStudy (anonymous):

yw

OpenStudy (goformit100):

Thanks @satellite73